OpenStudy (anonymous):

A bag contains 8 green marbles, 16 red marbles, 3 white marbles, and 3 black marbles. What is the probability of drawing a white marble out of the bag?

OpenStudy (anonymous):

Number of white marbles = 3 Total number of marbles in the bag = 30 Probability (that the marble drawn out is white ) = 3/30 = 1/10